Okay so if angle a and angle t are complementary and angle t is three times the measure of angle a what's the measure of angle a?
Answer by lwsshak3(11628)

Okay so if angle a and angle t are complementary and angle t is three times the measure of angle a what's the measure of angle a?
..
a + t=90º
t=3a
a + 3a=90º
4a=90º
a=22.5º
t=90-a=67.5º